OpenStudy (anonymous):

Fiona decides to join a gym for a minimum of one month to train for a triathlon. The gym charges a joining fee of $100 and a monthly fee of $38. If x represents the number of months that Fiona is a member of the gym, the equation below can be used to determine C, her total membership fee for that duration of time. 100 + 38x = C Fiona has allocated a maximum of $328 to spend on her gym membership. Which number line shows the possible number of months that Fiona can be a member of the gym?

OpenStudy (anonymous):

6 i think

OpenStudy (anonymous):

328-100=228/38=6

OpenStudy (anonymous):

is that 0-6 on a graph???

OpenStudy (anonymous):

I don't know I have troubles doing graphs

OpenStudy (anonymous):

me toooooo but thank you for trying to help